Find two consecutive integers such that four times the the larger exceeds three times the smaller by 23.
Answer by rapaljer(4671)

Let x = first (smaller)
x+ 1 = second (larger)

4(x+1) - 3(x) = 23
4x +4 - 3x = 23
x+4=23
x= 19
x+1 = 20

Check: 4(20)= 80
3(19) = 57

Difference = 23

R^2